OpenStudy (anonymous):

Solve for x lnx +ln(x-1) = 1

OpenStudy (anonymous):

Start by using both sides as exponents for e.

OpenStudy (anonymous):

ln a + ln b = ln ab

OpenStudy (anonymous):

e ^(a + b) - (e^a)(e^b). And e^(ln x) = x

OpenStudy (anonymous):

ln(x)(x-1) =1

OpenStudy (anonymous):

Is what I have so far good? Where do I go from here?

OpenStudy (anonymous):

e^[lnx +ln(x-1)] = e^1 = e. [e^(ln x)][e^(ln [x-1])] = x(x-1) = e = x^2 - x. Now, just solve quadratic.

OpenStudy (anonymous):

I tried solving it and I got x = 2.22 a,d x = -1.22 and I was told the negative could not be used so I used 2.22 and when I typed it in to my online math thing it said it was wrong

OpenStudy (anonymous):

and*

OpenStudy (anonymous):

and x > 1 because the original equation has an ln (x-1), so x-1 has to be positive.

OpenStudy (anonymous):

never mind I u=guess I wasn't putting in enough digits, it is correct, thanks!

OpenStudy (anonymous):

you're welcome

hero (hero):

Hey do this: \[lnx +ln(x-1) = 1\] Re-write the equation as follows: \[ln(x(x-1)) =1\] \[ln(x^2 - x) = 1\] Take the inverse ln of both sides to get \[x^2 - x = e\] Subtract e from both sides to get \[x^2 - x - e = 0\] Now use the quadratic formula where a = 1, b = -1, c = -e to solve for x